# CRYSP: a Total-Body PET based on cryogenic cesium iodide crystals

### v1

@@ IV-A Timestamp
The usual strategy is to approximate timestamp values by $T_{0}$ crossing time, whose precision is limited by timing resolution. A simple Time-to-Digital Converter (TDC) scheme can easily achieve resolutions below 100 ps with relatively slow clocks (20 ns period) [27] using affordable ADC converters (less than 9 bits). However, there is a significant effect related to signal slope behavior (time walk) that blurs effective timestamp resolution [28]. In order to reduce this effect a very low threshold can be employed. Therefore, false pulse detection probability would increase due to noise in the input stages of the front-end. Introducing control mechanisms to reject false pulse detection would increase hardware complexity. However, a different approach will be introduced in the development of this integrated front-end: different time values of the pulse waveform will be used to predict an enhanced timestamp value.
The proposed time-walk correction algorithm relies on several threshold crossing times of the rising slope of the pulse signal (fig. 9). The time intervals between these points along with the first (and lowest) threshold crossing point $T_{0}$, can be used to predict the true origin of the pulse signal:
@@ IV-B Pile-up processor
In a two-event pile-up scenario, only three key points are required to determine the energy of both events (see fig. 10). Two of these points, $A$ and $C$, represent the peak positions of the first and second events, respectively. The third point, $B$, corresponds to the start of the second event, which overlaps with the tail of the first event. In an event without pile-up, the peak magnitude is directly related to the event’s energy, as the signal shape consistently follows the same time distribution function. Consequently, the energy of the first event can be accurately determined from the amplitude at point $A$.
@@
Figure 10: Example of signal with pile-up (top) and its derivative (bottom), showing the latching times for $J_{1}$ and $J_{2}$.
This operation is an approximation to a deconvolution of the down slope effect of the first pulse, and can be computed offline after data acquisition. Only three data must be converted to digital and sent to the DAQ system: $A$ peak value, $B$ time position and $C-B$ difference (second event peak amplitude to be corrected offline). However, these data must be accurately sampled and latched timely enough to accomplish the ADC operations. The proposed algorithm can be implemented as a modified multiple peak detector (fig. 11) whose control signals rely on the aforementioned analog derivative computation. First of all, the capacitor $C_{A}$ stores the first peak and the $J_{1}$ control signal latches its value, which prevents it from losing peak information when the signal rises above it again. $J_{1}$ will be generated with the first change of the derivative signal from positive-to-negative values. At the same time, $J_{1}$ breaks $C_{B}$ charge path through $D_{1}$, so it turns into a falling peak detector along with $D_{2}$. When $J_{2}$ is activated by a negative-to-positive change in the derivative signal, $C_{B}$ gets latched with $B$ point value. In the second event, $C_{C}$ acts now as a peak detector along with $D_{1}$, then the second peak value gets stored in this capacitor until the maximum value is reached.
Given the timing characteristics of the detector signals and the method used to generate the peak detector control signals, there is a limit to the allowable time overlap between pile-up pulses. A conservative estimate on this limit gives a value of approximately 500 ns. The effect of the unresolved pile-up is quantified in detail in section VII-D.
@@
Figure 11: Modified multiple peak detector for pile-up processing. The $C_{A}$ capacitor stores the first peak, latched by $J_{1}$ upon the derivative signal’s positive-to-negative transition. $J_{1}$ disables $C_{B}$ charging path, turning it into a falling peak detector. $J_{2}$, triggered by a negative-to-positive transition, latches $C_{B}$ to point $B$, while $C_{C}$ detects the second peak with $D_{1}$, storing its maximum value.
V Modeling CRYSP
@@
A detailed microphysical simulation of CRYSP has been performed using the Geant4 simulation framework [13]. Two different applications have been implemented: one detailing a single monolithic crystal read out by a matrix of $8\times 8$ SiPMs of size $6\times 6$ mm2 and one with the baseline design of a full CRYSP1M scanner.
@@
The first application was used to study in detail the performances of monolithic crystals. The emission spectrum of CsI at cryogenic temperatures and the photon detection efficiency of the photosensor were specified according to fig. 5. Several possible wrapping materials (e.g., PTFE, ESR [30]) and coverage combinations were tested. The data obtained with the experimental apparatus of ref. [12] were used to validate the simulation. Optical photon generation was included in the application and the best agreement between data and Monte Carlo was obtained with the LUT optical model [31]. The results of the simulation were then used to develop the vertex reconstruction algorithm described in section VI.
@@
The second application includes the full geometry of CRYSP1M, as shown in fig. 12, and was used for the performance evaluation detailed in sections VII and VIII. Positrons inside the phantoms were generated with a kinetic energy corresponding to the 18F decay spectrum. However, simulation of optical photons is too computationally intense for a full detector. Thus, the energy deposition in the crystals is smeared with a Gaussian FWHM of 6.3%, corresponding to the experimental energy resolution obtained in ref. [12]. The light time emission profile and CTR are applied separately at analysis level.
Figure 12: Geometry simulated with Geant4 of the CRYSP1M detector with a modified Jaszczak phantom placed at the center of the FOV. An electron-positron annihilation inside the phantom produces two gamma rays which are detected by scintillating crystals. Optical photons are drawn here but not simulated in the full CRYSP1M geometry.
VI Reconstruction of interaction vertex in CRYSP
VI-A Convolutional neural network for vertex reconstruction
@@
In a pixelated detector, the reconstruction of the interaction vertex (iV) usually requires a single pixel fired in the detector unit, typically built as an array of several pixels (e.g., in the Quadra scanner, a mini block is made of $5\times 5$ pixels, and each pixel has transverse dimensions of $3.2\times 3.2~{}$mm2, thus a detector unit is a matrix of 25 pixels with dimensions $16\times 16~{}$mm2). Imposing that the energy deposited in the pixel is near 511 keV selects events at or very near the photoelectric peak. The transverse position of the iV is known with a resolution of the order $\mathrm{FWHM}_{t}=2\sqrt{2\ln 2}\times t/\sqrt{12}$, where $t$  denotes the transverse dimension of the pixel. For example, in the case of the Quadra, $t=3.2$ mm and thus $\mathrm{FWHM}_{t}=2.1$ mm. The depth of interaction (d.o.i.) is known with a resolution $\mathrm{FWHM}_{doi}=2\sqrt{2\ln 2}\times l/\sqrt{12}$, where $l$ is the length of the crystal. Considering again the example of the Quadra, $l=20$ mm and $\mathrm{FWHM}_{doi}=13.3$ mm. Thus, while the small size of the pixels guarantees excellent transverse resolution (at the cost of a large number of channels), the d.o.i resolution is typically poor, even for dense crystals, such as LYSO or BGO, since one requires typically a minimum of 2 radiation lengths ($X_{0}$) for acceptable sensitivity. An additional problem, whose impact becomes increasingly important as the AFOV of the scanner increases, is that of parallax. Gammas entering the detectors at very large angle often do not deposit their energy in the first pixel they encounter on their trajectory, but one or two pixels further. Parallax results in large errors in the reconstruction of the iV. Additionally, gammas can scatter in one pixel, depositing a fraction of their energy, before being fully absorbed in a neighboring pixel. Such events are typically rejected, negatively impacting the sensitivity.
@@
CsI has a longer radiation length ($X_{0}\sim 18.5$ mm) and smaller photoelectric fraction ($\sim 20$%) than the dense, high $Z_{eff}$ scintillators (BGO, LYSO), usually deployed in modern PET scanners. In a pixelated detector this would translate to a larger error in the d.o.i, as well as in a smaller fraction of selected events.
@@
CRYSP, instead, proposes the use of monolithic crystals, of relatively larger dimensions, read out by an array of large-area SiPMs (see section III). One immediate advantage of such a choice is to reduce the number of readout channels by approximately a factor of 4 (from $3\times 3$ mm2 to $6\times 6$ mm2 SiPMs). On the other hand, the simple algorithm used for pixelated systems cannot be applied in this case, since all the 64 SiPMs reading out the crystal have signal when a gamma interacts in the crystal.
@@
The adoption of monolithic crystals for PET and their potential regarding iV reconstruction is not a novel topic [32, 33]. However, obtaining a resolution comparable with the transverse dimensions has been historically challenging and classical algorithms fare relatively poorly [34]. Algorithms based on multilayer perceptrons have also been implemented, providing similar results [35, 36]. Here, instead, we show how excellent resolution in all three coordinates of the iV can be achieved using CNNs, which have been widely adopted for image recognition tasks with extraordinary success [37, 38].
@@
Furthermore, the use of CNNs allows to include in the sample events with a first Compton scatter followed by photoelectric absorption, which deposit the full 511 keV in the same crystal. This is a much larger fraction than pure photoelectric events, up to 60% of the total events, and is also more significant in CsI than in higher $Z_{eff}$ materials such as LYSO and BGO. Most of these events deposit their energy in two or more clusters in the crystal, but the two most energetic clusters add up to a very large fraction of the total energy deposited, and thus determine the shape of the image. Notably, the first true interaction vertex (e.g., the point at which the gamma truly interacts in the crystal) coincides most of the time with the first apparent interaction vertex (e.g., the vertex closer to the entry point of the gamma), thus resulting in unbiased, gaussianly-distributed residuals between the true interaction vertex of the gamma and the first apparent vertex predicted by our CNN-based reconstruction algorithm, as shown in fig. 14.
@@
Our CNN is, therefore, trained to search for two iVs $(x_{1},y_{1},z_{1},x_{2},y_{2},z_{2})$ where the index 1 is assigned to the first apparent interaction vertex. Training the CNN to find two vertices rather than one results in better determination of the position (since, in practice, an algorithm trained to find a single vertex determines the barycenter of the interaction which may be several mm away from the true vertex), and assigning the iV to the first apparent vertex is a good approximation to the true vertex. The full network is made of three consecutive convolutional blocks (Conv2D + BatchNorm + LeakyReLU + MaxPooling2D) followed by a Flatten, a Dropout, and a Dense layer yielding 6 outputs $(x_{1},y_{1},z_{1},x_{2},y_{2},z_{2})$, as shown in fig. 13.
Figure 13: Model of the neural network used to predict two iVs in the crystal, taking as input a map of the photoelectrons detected by each SiPM. The convolutional block is repeated three times and then fed to a dense layer, after passing through a flatten and a dropout layer.
@@
The slightly larger error in the determination of the transverse coordinates, compared with pixelated PET scanners, is compensated by a much smaller error on the d.o.i. Notice, also, that the resolution in the case of a monolithic crystal stays roughly constant with the transaxial angle, since the determination of the iV is not affected by parallax (see sec. VII-B). For the spatial resolution and image reconstruction studies described in the following sections, the position of the first interaction of the gamma within the crystal is smeared with the Gaussian distributions of fig. 14.
@@ VII-A Sensitivity
The procedure was repeated in two different configurations, one with the phantom placed at the center of the FOV and one with a radial offset of 10 cm, obtaining $R_{\mathrm{0~{}cm}}=120\pm 5$ kcps/MBq and $R_{\mathrm{10~{}cm}}=128\pm 5$ kcps/MBq, respectively. Although these values are lower than those reported for LYSO-based TBPETs (174 kcps/MBq for the uEXPLORER [40] and 176 kcps/MBq for the Quadra [4]), the selected events exhibit less Compton contamination due to superior energy resolution. This results in reconstructed images of comparable quality at parity of number of events (see sec. VIII).
The attenuation-corrected sensitivity as a function of the axial position for both radial offsets is reported in fig. 15.
VII-B Spatial resolution
Spatial resolution of the scanner has been quantified according to the NEMA standard by simulating a point-like positron source placed at six different positions: in the axial direction at the center of the AFOV and three-eights of the AFOV from the center and in the transverse direction at 1 cm, 10 cm, and 20 cm from the center. The resolution is measured in all three directions: radially, axially, and tangentially. Table III shows the full width at half-maximum (FWHM) and full width at tenth-maximum (FWTM) amplitudes for the six positions and for maximum likelihood expectation maximization (MLEM) with 10 iterations.
TABLE III: FWHM and FWTM in all three directions for point-like sources at six different locations, obtained with the MLEM algorithm.
Axial position (mm) Radial position (mm) Full width at half-maximum (mm) Full width at tenth maximum (mm)

### v2

@@ 2.3.1 Timestamp
The usual strategy is to approximate timestamp values by $T_{0}$ crossing time, whose precision is limited by timing resolution. A simple Time-to-Digital Converter (TDC) scheme can easily achieve resolutions below 100 ps with relatively slow clocks (20 ns period) [kalisz2003review] using affordable ADC converters (less than 9 bits). However, there is a significant effect related to signal slope behavior (time walk) that blurs effective timestamp resolution [time_walk]. In order to reduce this effect a very low threshold can be employed. Therefore, false pulse detection probability would increase due to noise in the input stages of the front-end. Introducing control mechanisms to reject false pulse detection would increase hardware complexity. However, a different approach will be introduced in the development of this integrated front-end: different time values of the pulse waveform will be used to predict an enhanced timestamp value.
The proposed time-walk correction algorithm relies on several threshold crossing times of the rising slope of the pulse signal (see fig. 6). The time intervals between these points along with the first (and lowest) threshold crossing point $T_{0}$, can be used to predict the true origin of the pulse signal:
@@ 2.3.2 Pile-up processor
In a two-event pile-up scenario, only three key points are required to determine the energy of both events (see fig. 7a). Two of these points, $A$ and $C$, represent the peak positions of the first and second events, respectively. The third point, $B$, corresponds to the start of the second event, which overlaps with the tail of the first event. In an event without pile-up, the peak magnitude is directly related to the event’s energy, as the signal shape consistently follows the same time distribution function. Consequently, the energy of the first event can be accurately determined from the amplitude at point $A$.
@@
Figure 7: (a) Example of signal with pile-up (top) and its derivative (bottom), showing the latching times for $J_{1}$ and $J_{2}$. (b) Modified multiple peak detector for pile-up processing. The $C_{A}$ capacitor stores the first peak, latched by $J_{1}$ upon the derivative signal’s positive-to-negative transition. $J_{1}$ disables $C_{B}$ charging path, turning it into a falling peak detector. $J_{2}$, triggered by a negative-to-positive transition, latches $C_{B}$ to point $B$, while $C_{C}$ detects the second peak with $D_{1}$, storing its maximum value.
This operation is an approximation to a deconvolution of the down slope effect of the first pulse, and can be computed offline after data acquisition. Only three data must be converted to digital and sent to the DAQ system: $A$ peak value, $B$ time position and CBC-B difference (second event peak amplitude to be corrected offline). However, these data must be accurately sampled and latched timely enough to accomplish the ADC operations. The proposed algorithm can be implemented as a modified multiple peak detector (fig. 7b) whose control signals rely on the aforementioned analog derivative computation. First of all, the capacitor $C_{A}$ stores the first peak and the $J_{1}$ control signal latches its value, which prevents it from losing peak information when the signal rises above it again. $J_{1}$ will be generated with the first change of the derivative signal from positive-to-negative values. At the same time, $J_{1}$ breaks $C_{B}$ charge path through $D_{1}$, so it turns into a falling peak detector along with $D_{2}$. When $J_{2}$ is activated by a negative-to-positive change in the derivative signal, $C_{B}$ gets latched with $B$ point value. In the second event, $C_{C}$ acts now as a peak detector along with $D_{1}$, then the second peak value gets stored in this capacitor until the maximum value is reached.
Given the timing characteristics of the detector signals and the method used to generate the peak detector control signals, there is a limit to the allowable time overlap between pile-up pulses. A conservative estimate on this limit gives a value of approximately 500 ns. The effect of the unresolved pile-up is quantified in detail in section 3.4.
2.4 Modeling CRYSP
@@
A detailed microphysical simulation of CRYSP has been performed using the Geant4 simulation framework [GEANT4:2002zbu]. Two different applications have been implemented: one detailing a single monolithic crystal read out by a matrix of $8\times 8$ SiPMs of size $6\times 6$ mm2 and one with the baseline design of a full CRYSP1M scanner.
@@
The first application was used to study in detail the performances of monolithic crystals. The emission spectrum of CsI at cryogenic temperatures and the photon detection efficiency of the photosensor were set according to specifications [Hamamatsu]. Several possible wrapping materials (e.g., PTFE, ESR [loignon2017reflectivity]) and coverage combinations were tested. The data obtained with the experimental apparatus of ref. [Soleti:2024flw] were used to validate the simulation. Optical photon generation was included in the application and the best agreement between data and Monte Carlo was obtained with the LUT optical model [5485130]. The results of the simulation were then used to develop the vertex reconstruction algorithm described in section 2.5.
@@
The second application includes the full geometry of CRYSP1M and was used for the performance evaluation detailed in section 3. Positrons inside the phantoms were generated with a kinetic energy corresponding to the 18F decay spectrum. However, simulation of optical photons is too computationally intense for a full detector. Thus, the energy deposition in the crystals is smeared with a Gaussian FWHM of 6.3%, corresponding to the experimental energy resolution obtained in ref. [Soleti:2024flw]. The light‐time emission profile and coincidence time resolution (CTR) were applied separately at the analysis stage. A scintillation decay constant of 800 ns and a CTR of 1.5 ns FWHM were adopted, consistent with the experimental results reported in ref.  [Soleti:2024flw]. The decay time was implemented by time-distributing the detected photoelectrons according to an exponential emission profile with the corresponding decay constant, thereby reproducing the temporal response of the photosensor to the scintillation pulse. Dead-time effects were modeled through the integration and pile-up processing described in section 2.3.2. Interactions occurring within 500 ns were classified as unresolved pile-up and treated as a single event, while pulses separated by more than 500 ns were assumed to be resolvable by the electronics and to preserve the nominal energy resolution.
The acquisition parameters used in all performance studies are consistent with those expected for the final detector operation and are summarized in Table 2. In particular, events were selected within an energy window of 465–555 keV (corresponding to $3\sigma$ around 511 keV) and a coincidence time window of 4.5 ns. These settings were adopted for both sensitivity and imaging simulations to allow a fair comparison across all performance metrics.
In order to provide a consistent benchmark, we also simulated a reference LYSO-based PET scanner under the same Monte Carlo framework and acquisition settings used for CRYSP. The geometry and axial coverage were matched to the CRYSP configuration, with detector modules composed of pixelated LYSO crystals ($3\times 3\times 22$ mm3) read out by SiPMs. The intrinsic parameters were set to 10% energy resolution, 40 ns decay time, an energy window of 435–585 keV (also here corresponding to $3\sigma$ around 511 keV), and a 350 ps CTR, which lies midway between the performances reported for the Biograph Vision Quadra (230 ps) [prenosil2022performance] and the uEXPLORER (505 ps) [badawi2019first], which enables TOF reconstruction and is consistent with reported performances of state-of-the-art TBPET systems such as the Biograph Vision Quadra [prenosil2022performance] and uEXPLORER [badawi2019first]. This virtual LYSO scanner serves as a physically consistent reference for comparing sensitivity, NECR, and spatial resolution results, complementing the image-reconstruction comparison, as shown in section 3.
2.5 Reconstruction of interaction vertex in CRYSP
In a pixelated detector, the reconstruction of the interaction vertex (iV) usually requires a single pixel fired in the detector unit, typically built as an array of several pixels (e.g., in the Quadra scanner, a mini block is made of $5\times 5$ pixels, and each pixel has transverse dimensions of $3.2\times 3.2$~mm2, thus a detector unit is a matrix of 25 pixels with dimensions $16\times 16$~mm2). Imposing that the energy deposited in the pixel is near 511 keV selects events at or very near the photoelectric peak.
The intrinsic spatial resolution in such a configuration is mainly determined by the crystal pitch and thickness. The transverse position of the interaction vertex (iV) can be approximated, in first order, by $\mathrm{FWHM}_{t}=2\sqrt{2\ln 2}\times t/\sqrt{12}$, where $t$ denotes the transverse dimension of the pixel. For instance, in the Biograph Vision Quadra scanner, $t=3.2$ mm resulting in $\mathrm{FWHM}_{t}=2.1$ mm. Similarly, the d.o.i. resolution can be estimated as $\mathrm{FWHM}_{doi}=2\sqrt{2\ln 2}\times l/\sqrt{12}$, where $l$ is the length of the crystal. Considering again the example of the Quadra, $l=20$ mm yields $\mathrm{FWHM}_{doi}=13.3$ mm. These expressions represent simple geometrical estimates assuming a uniform interaction distribution and are used as analytical proxies for the intrinsic sampling limits of pixelated PET detectors.
@@ 2.5 Reconstruction of interaction vertex in CRYSP
Thus, while the small size of the pixels guarantees excellent transverse resolution (at the cost of a large number of channels), the d.o.i resolution is typically poor, even for dense crystals, such as LYSO or BGO, since one requires typically a minimum of 2 radiation lengths ($X_{0}$) for acceptable sensitivity. An additional problem, whose impact becomes increasingly important as the AFOV of the scanner increases, is that of parallax. Gammas entering the detectors at very large angle often do not deposit their energy in the first pixel they encounter on their trajectory, but one or two pixels further. Parallax results in large errors in the reconstruction of the iV. Additionally, gammas can scatter in one pixel, depositing a fraction of their energy, before being fully absorbed in a neighboring pixel. Such events are typically rejected, negatively impacting the sensitivity.
@@
CsI has a longer radiation length ($X_{0}\sim 18.5$ mm) and smaller photoelectric fraction than the dense, high $Z_{eff}$ scintillators (BGO, LYSO), usually deployed in modern PET scanners. In a pixelated detector this would translate to a larger error in the d.o.i, as well as in a smaller fraction of selected events.
@@
CRYSP, instead, proposes the use of monolithic crystals, of relatively larger dimensions, read out by an array of large-area SiPMs (see section 2.2). The simple algorithm used for pixelated systems cannot be applied in this case, since all the 64 SiPMs reading out the crystal have signal when a gamma interacts in the crystal.
@@
The adoption of monolithic crystals for PET and their potential regarding iV reconstruction is not a novel topic [gonzalez2019novel, gonzalez2021evolution]. However, obtaining a resolution comparable with the transverse dimensions has been historically challenging and classical algorithms fare relatively poorly [li2008high]. Algorithms based on multilayer perceptrons have also been implemented, providing similar results [freire2022position, iborra2019ensemble]. Here, instead, we show how excellent resolution in all three coordinates of the iV can be achieved using CNNs, which have been widely adopted for image recognition tasks with extraordinary success [li2021survey, yamashita2018convolutional].
@@
Furthermore, the use of CNNs allows to include in the sample events with a first Compton scatter followed by photoelectric absorption, which deposit the full 511 keV in the same crystal. This is a much larger fraction than pure photoelectric events, up to 60% of the total events, and is also more significant in CsI than in higher $Z_{eff}$ materials such as LYSO and BGO. Most of these events deposit their energy in two or more clusters in the crystal, but the two most energetic clusters add up to a very large fraction of the total energy deposited, and thus determine the shape of the image. Notably, the first true interaction vertex (e.g., the point at which the gamma truly interacts in the crystal) coincides most of the time with the first apparent interaction vertex (e.g., the vertex closer to the entry point of the gamma), thus resulting in unbiased, gaussian-distributed residuals between the true interaction vertex of the gamma and the first apparent vertex predicted by our CNN-based reconstruction algorithm, as shown in fig. 8.
@@
Our CNN is, therefore, trained to search for two iVs $(x_{1},y_{1},z_{1},x_{2},y_{2},z_{2})$ where the index 1 is assigned to the first apparent interaction vertex. Training the CNN to find two vertices rather than one results in better determination of the position (since, in practice, an algorithm trained to find a single vertex determines the barycenter of the interaction which may be several mm away from the true vertex), and assigning the iV to the first apparent vertex is a good approximation to the true vertex. The full network is made of three consecutive convolutional blocks (Conv2D + BatchNorm + LeakyReLU + MaxPooling2D) followed by a Flatten, a Dropout, and a Dense layer yielding 6 outputs $(x_{1},y_{1},z_{1},x_{2},y_{2},z_{2})$.
@@
The slightly larger error in the determination of the transverse coordinates, compared with pixelated PET scanners, is compensated by a much smaller error on the d.o.i. Notice, also, that the resolution in the case of a monolithic crystal stays roughly constant with the transaxial angle, since the determination of the iV is not affected by parallax (see section 3.2).
For the spatial resolution and image reconstruction studies of CRYSP described in the following sections, the positions of the first interaction of the two gammas within the crystals are smeared with the two Gaussian distributions of fig. 8 and are then used directly to define the listmode LORs. Although a mild depth dependence of the spatial resolution exists (0.2–0.5 mm), in a monolithic detector the d.o.i. is not known a priori and must be estimated from the light distribution. Thus, as the reconstruction algorithm does not have access to the true depth, a depth-dependent intrinsic smearing cannot be applied in a realistic system-level simulation. In the pixelated LYSO-based PET system, where no dedicated d.o.i. encoding is assumed, the interaction position is assigned to the geometric position of the crystal.
@@ 3.1 Sensitivity
The procedure was repeated in two different configurations, one with the phantom placed at the center of the FOV and one with a radial offset of 10 cm, obtaining $R_{\mathrm{0~cm}}=120\pm 5$ kcps/MBq and $R_{\mathrm{10~cm}}=128\pm 5$ kcps/MBq, respectively. Although these values are lower than those reported for LYSO-based TBPETs (174 kcps/MBq for the uEXPLORER [spencer2021performance] and 176 kcps/MBq for the Quadra [prenosil2022performance]), the selected events exhibit less Compton contamination due to superior energy resolution, as detailed in section 3.3. The attenuation-corrected sensitivity as a function of the axial position for both radial offsets is reported in fig. 9. This results in reconstructed images of comparable quality at parity of number of events (see section 3.5). For reference, our simulated LYSO-based PET scanner yields a sensitivity of $(177\pm 6)$ kcps/MBq, in excellent agreement with the reported Quadra performance of 175.3 kcps/MBq [prenosil2022performance], further supporting the validity of our Monte Carlo model.
3.2 Spatial resolution
Spatial resolution of the scanner has been quantified according to the NEMA standard by simulating a point-like positron source placed at six different positions: in the axial direction at the center of the AFOV and three-eights of the AFOV from the center and in the transverse direction at 1 cm, 10 cm, and 20 cm from the center. The resolution is measured for both the CRYSP and the reference LYSO-based PET system in all three directions: radially, axially, and tangentially. Table 3 shows the full width at half-maximum (FWHM) and full width at tenth-maximum (FWTM) amplitudes for the six positions obtained with the filtered backprojection algorithm. Notably, the CRYSP performance becomes significantly superior to that of the LYSO-based PET at large radial offsets and off-center axial positions, where parallax effects in pixelated detectors degrade spatial resolution.
Table 3: FWHM and FWTM in all three directions for point-like sources at six different locations in CRYSP and in the LYSO-based reference PET, obtained with the filtered backprojection algorithm.
System Axial position (mm) Radial offset (mm) FWHM (mm) FWTM (mm)